OpenStudy (anonymous):

Can someone help me for a medal and fan?!?! An equation is shown below: 9x - y = -2

OpenStudy (anonymous):

@aliboo997 @aum

OpenStudy (anonymous):

can you help me in return

OpenStudy (anonymous):

It really depends on the problem. I'll try though

OpenStudy (anonymous):

what do you want to find x and y

OpenStudy (anonymous):

Well really what I want is to know how to do this

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

add o into x to get y and add o into y to get x

OpenStudy (anonymous):

That doesn't make sense

OpenStudy (aum):

The problem sates: "An equation is shown below: 9x - y = -2" But there is no question. So what are they asking you to do?

OpenStudy (anonymous):

Oh, my bad. Hang on

OpenStudy (anonymous):

Sorry thought I added it. Here it is

OpenStudy (anonymous):

Explain how you will show all of the solutions that satisfy this equation.

OpenStudy (anonymous):

If you want, I can combine it all to make it easier on you

OpenStudy (aum):

9x - y = -2 is the equation of a straight line. ALL points ON the line are solutions to the equation. But if we don't want to graph it but wish to answer this algebraically, then let us solve for y first: Add y to both sides: 9x = y - 2 add 2: 9x + 2 = y or y = 9x + 2. We can pick ANY value for x, substitute it into the equation above and find the corresponding y. y = 9x + 2 x y 0 2 1 11 2 20 -1 -7 -2 -16 ....

OpenStudy (aum):

So the solutions are: (-1, -7), (-2, -16), (0, 2), (1, 11), (2, 20), .......

OpenStudy (anonymous):

Okay, gotcha. Sorry for the delay, somebody knocked. I can finish it from here. Thank you!